### 24 Mar 2026 v1

Abstract
Integrated opto-electronic devices have the potential to revolutionize information processing, with substantial increase in compute speed, seamless information transfer and reduction of energy consumption. A key missing unit for the successful implementation of compact functional devices are nanometer scale modular and tunable light sources. Monotonically grown semiconducting nanowire lasers (NWLs) fills this gap. However, NWLs operation improvement and optimization require the characterization of their near-field and its dynamics at the nanometer scale, which is hindered due to the light diffraction limit. Here we show how synchronous electron near-field and photon far-field time-resolved spectroscopies surpass this limitation and map a NWLs near-field with nanometer and sub-picoseconds temporal resolution. We quantitatively measured the evolution of the absolute number of stimulated photons in the NWL cavity, measuring that up to 4x105 are present simultaneously in the cavity. We mapped the lasing cavity mode’s near-field, showing that both whispering gallery and Fabry-Perot modes can participate in the lasing. Our results demonstrate how the near-field of a NWL under operation evolves in the sub-picoseconds and the nanometer scales. We anticipate that a direct observation of the near-field will help to elucidate the influence of materials heterogeneities (defects, chemical changes, contaminants, interface roughness, strain) in NWL operation.
keywords:
UTEM, NanoLaser, PINEM, light-matter interaction, out-of-equilibrium spectroscopySemiconductors nanowire lasers (NWLs) are one of the most promising new class of nanometer scale light sources, with sub-wavelength applications in bio-sensing and on-chip integrated photonics. Indeed, their monolithic growth insures both a high material quality and an economical viability [ellis_nanolasers_2024, behzadirad_scalable_2018, damilano_top-down_2019]. In addition, their confined cavity mode guarantee a low lasing threshold [gradecak_gan_2005, couteau_nanowire_2015] and ultrafast dynamics [roder_ultrafast_2015]. Following the initial experimental demonstration of a ZnO nanowire laser in 2001 [huang_room-temperature_2001], NWLs exploiting II-VI, III-V [johnson_single_2002, saxena_optically_2013], and various other materials [yang_room_2017] have been developed. The abundance of semiconductor gain materials enables a significant tunability for specific applications (wavelength, material compatibility, etc.). One of their most interesting characteristic is that the cavity properties are essentially determined by the geometry of the nanowire, due to the significant refractive index difference between vacuum and semiconductor medium [piprek_gan_2007]. This means that the lasing wavelength of a NWL is mainly dictated by the cavity length, a design parameter which can be easily tuned [piprek_gan_2007, zhuang_multicolor_2019]. However, despite the dramatic influence of these sub-wavelength parameters, the optical characteristics of nanolasers (e.g. lasing threshold, power density, mono-chromaticity, and lasing dynamics [roder_ultrafast_2015]) have up to now been investigated in the far-field region. There is thus a strong incentive to push the study of NWL in the near-field i.e. accessing the evanescent field with a subwavelength resolution. This was for example achieved on photonic crystal micro-lasers using scanning near-field optical microscopy (SNOM) [vo_near-field_2010], where an efficient mapping of the luminescence properties was reported. However, it was also shown that the near-field probe has a strong influence on the measured properties [le_gac_tuning_2009], hindering the measurement of intrinsic properties. In this article, we report the first synchronized measurement of far-field and spatially resolved sub-picoseconds temporal dynamics of the optical near-field of NWLs during lasing operation. To this aim, we use a combination of micro photo-luminescence spectroscopy (PL) and photon induced near-field electron microscopy (PINEM) [barwick_photon-induced_2009] implemented in a custom-built ultrafast transmission electron microscope (UTEM). In the following, we will demonstrate: (1) the quantitative measurement of the absolute number of stimulated photons inside the NWL cavity as a function of time with sub-picoseconds resolution, and (2) the mapping of the lasing cavity mode’s near-field with nanometer resolution.
1 PINEM on Nanowire laser
We studied GaN nanowires with 50 nm InGaN inclusions every 150 nm (Fig. 1b). The nanowires are fabricated using a top-down technique (see SI) [damilano_top-down_2019]. The UTEM used in this experiment is a modified Cold-FEG HF3300 Hitachi [houdellier_development_2018], wherein a 250 fs ultrashort laser pulsed at 2 MHz is directed to both the electron gun to generate a 400 fs pulsed electron beam, and to the sample via a parabolic mirror. The latter enables the efficient excitation of a single NWL with a spot size of approximately 10 , as well as the collection of its emission spectrum (PL, see Fig. 1c). Crucially, the electron energy and the PL spectra are recorded simultaneously, so that the measurements of near- and far-field of the NWLs are always correlated. It is particularly important as the stimulated emission of the NWL is marked by the emergence of sharp PL peaks [couteau_nanowire_2015] at the cavity modes frequencies (Fig. 1c). The power input vs. intensity output of a typical NWL - in the condition of illumination of the experiment and in a standard micro-photoluminescence (PL) experiment - are shown in SI (Fig.6 and Fig. 13).
PINEM is a pump-probe technique where a focused pulsed electron beam probes the z component of the near-field induced by a pulsed laser on a nano-structure (see Fig. 1a). Before this work, in PINEM, the optical modes have always been resonantly excited by the pulsed laser and did not involved light absorption processes. Indeed here, the optical pumping produces electrons and holes relaxing and providing gain, leading to laser emission in a specific electromagnetic mode of the nanowire. When passing through the excited near-field of an object, the electron exchanges quanta of energy with the optical field, resulting in an electron energy spectrum with sidebands spaced by the probed photon energy (see Fig. 1d). The amplitude of the sideband - corresponding to the electron having exchanged with the field - is noted with . Remarkably, all the amplitudes are uniquely determined by a single quantity - the electron-light coupling constant - which can be expressed as [di_giulio_probing_2019]:
Here, denotes the electron-photon scattering amplitude - a mode property independent of the laser excitation - which can be measured by electron energy-loss spectroscopy (EELS) [nelayah_mapping_2007]. Therefore, the dynamical properties of the NWL is contained in and can be reconstructed by varying the delay between the laser pump and the electron probe. The number of photons in the cavity as a function of time is evaluated from .
2 Results
2.1 Temporal dynamics
We started by focusing the electron beam at a fixed position on the side of the NWL in an aloof geometry and recorded electron energy spectra for each delay between the pump laser pulse and the electron probe pulse (Fig. 1b), constructing what is called delay scans. These were recorded for a pumping below (P=0.25 mW) and above (P=0.66 mW) the lasing threshold (Fig. 2). For negative delays - corresponding to the probe impinging on the sample before the pump - the electron energy remains unchanged, and only the so-called zero-loss peak (ZLP) is observed, as expected.
Below the lasing threshold (Fig. 2a), the delay scan displays several energy-exchange sidebands around . Each sideband has a time-energy correlation due to the electron pulse energy-chirp [park_chirped_2012] of 0.89 ps/eV. This chirp is a consequence of the Coulomb repulsion between electrons within the electron pulse [haindl_coulomb-correlated_2023, feist_ultrafast_2017]. The Coulomb interaction increases the pulse width (from 400 fs to 800 fs), and is taken into account in our models [park_chirped_2012]. The energy spacing between each sideband is constant as a function of and measured at approximately 3.6 eV. This energy matches the pump laser wavelength, confirming that the electron interacts with the direct scattering of the pump on the surface of the nanowire.
The same direct scattering phenomenon is observed when the NWL is pumped above threshold (Fig. 2b). However, a second set of sidebands also appears at ps. They are separated by approximately 3.3 eV which corresponds to the energy of the stimulated emission visible in the PL spectrum (Fig. 1c). This second interaction can thus be identified to be the interaction between the electron beam and the stimulated near-field in the NWL cavity. While several cavity modes are visible on the PL spectra, the spectral resolution ( 1.0 eV) of our PINEM setup prevents us from separating them on the electron energy spectra. This is primarily due to the energy dispersion of the electron source used in this experiment.
To confirm our observation, we have simulated the delay scan expected for a lasing NWL with the electron chirp condition of our experiment (Fig. 2c and in SI Fig. 10). Our simulation and experiment display a remarkable agreement when the NWL output laser pulse width is modeled with an exponential decay of characteristic time of ps and an onset time ps. While the latter is mainly determined by the carriers thermalization, the former is known to be influenced by the cavity geometry and modes [roder_ultrafast_2015]. This short lifetime is in agreement with the expected ultrafast dynamic of nanolasers [johnson_ultrafast_2004, daskalakis_ultrafast_2018].
2.2 Photon population
We attribute the first set of electron energy sidebands ( ps) to the pump laser light directly scattered at the surface of the nanowire, and the second set of sidebands to the stimulated near-field of the NWLs during lasing. We will now extract the electron-light coupling constant from the sidebands amplitude , starting from the first set of sidebands. Since they correspond to the interaction between an electron pulse (400 fs) which is longer than the light pulse (250 fs), the scattering can be considered incoherent (each electron of the wavepacket experience a different field). In this case, the sidebands are modeled by a Poisson distribution [park_photon-induced_2010]. The resulting fit of is shown on Fig. 3a for three different pump powers. This approximation does not apply to the second set of sidebands, where the optical near-field has a long dynamics compared to the electron wavepacket duration. The resulting slight increase in the coherence of the interaction was modeled as in [harvey_probing_2020] and the resulting is extracted for the same three different pump powers.
The coupling constant variations are attributed to the photon population dynamics in the cavity (equation (1)). To obtain a quantitative estimate of the photon population, we have measured the electron-photon coupling constant of the same NWL using EELS on a 200 keV monochromated STEM (NION Hermes). The EELS spectrum taken at the same position as the PINEM delay scan is shown on Fig. 3c, a peak at 3.3 eV is clearly visible just below the band gap (marked by a red arrow). More peaks are visible at lower energy corresponding to other cavity modes but not coupled to the bath of spontaneous photons and therefore unavailable for lasing. From this experiment, we have estimated a value of and deduced the photons population dynamic in the cavity for each pump powers (Fig 3-b).
As expected [lorke_theory_2013, weng_electronhole_2021], we observe a clear decrease of the NWL rising time as a function of the pump power (Fig. 3). To further quantify this observation, the extracted are fitted with an exponential decay convoluted with , to take the electron chirp into account. The onset time shifts from 0.85 ps at P = 0.66 mW to 0.53 ps at P = 1.0 mW while the NWL pulse width is about 0.6 ps varying from 0.5 ps to 0.7 ps. Moreover, the cavity modes display a slight energy shift with increasing pump power (see the PL spectra in Fig. 3d), changing the ratio between modes, showing that it will be interesting to probe the different modes dynamic with PINEM individually. While our PINEM setup cannot spectrally resolve individual modes, PINEM in state-of-the-art monochromated TEM should reveal this inter-mode dynamics [castioni_nanosecond_2025]. Due to this lack of spectral resolution we cannot ascertain which of the modes we are actually probing or if all of them are participating to the near-field at this specific electron beam position (see discussion on near-field mapping).
2.3 Cavity mode imaging
So far, we have discussed the temporal dynamics of the photon inside the cavity, and will now turn to the spatial mapping of the near field of the lasing mode, with sub-wavelength resolution. Indeed, due to their large diameter (between 400 and 500 nm), the NWLs sustains two types of cavity excitation: Fabry-Perot modes (FPM) and whispering cavity modes (WGM) [coulon_gan_2012]. The far-field emission pattern is different for both kind of modes and are observed in PL experiments (see SI).
In order to characterize the near-field profiles of the bare cavity modes (i.e. in absence of external laser pumping), we performed an STEM-EELS hyperspectral imaging experiment on the same monochromated Nion Hermes machine - a technique which had already demonstrated its efficiency in probing such excitations [du_electron_2022, auad_unveiling_2022]. Fig. 4a shown a typical EELS study of these NWL. The spectrum taken at the side of the NWL exhibits clear peaks that are not visible at the top of the nanowire. We identified the WGM with the clear peaks present in the side while on the top of the NWL only a large number of FPM that almost seems continuous is visible [du_electron_2022]. Fitting the EELS spectrum around 3.3 eV we can separate the contribution to the WGM (amplitude of the peak) and to the Fabry Perrot (Amplitude of the continuous background). The different spatial variations of the cavity profile for FPW and WGM are clearly visible and shown on the Fig. 4a for E=3.3 eV.
We can now image the near-field of the NWL when illuminated by the pump laser above the lasing threshold. To do so, we perform the same PINEM experiment as before albeit here, the pump-probe delay is fixed while the impact parameter of the electron beam is raster-scanned over the nanowire. The resulting -maps for two delays - ps (i.e. direct scattering of the pump laser on the nanowire) and ps (lasing NWL cavity) - are shown on Fig. 4b. The second delay was chosen so because the pump power being slightly above the lasing threshold ( mW), the onset time is of about 2 ps after optical excitation (see SI).
For the first case (), the intensity of the near-field depends on the geometry of the different interfaces, on the aluminum layer below the nanolaser as well as on the angle between the incident light and the sample due to the parabolic mirror [meuret_photon-induced_2024]. However, for the second case showed ( ps), the induced near-field is drastically different and is clearly localized around two InGaN inclusions, spreading over a distance of 200 nm. By comparison with the EELS experiment (Fig. 4a), we can attribute this near-field profile to a WGM. Nevertheless, other cavity modes can be involved in the lasing, as shown on another NWL (NWL2) Fig. 4c where the near-field display clear features of a FPM - with two lobes located at the end facet. It is also interesting to note here that Fig. 4c shows a neighboring nanolaser (NWL3) that does not lase. PINEM allow us to separate the behavior of two very closely located NWL.
3 Conclusion
In conclusion, the spatial and temporal distribution of lasing modes in GaN NWL was mapped with few nanometer and sub-picosecond resolutions using PINEM. Coupled to monochromated EELS, this method revealed the time-dependent photon population within the cavity - without a priori knowledge on the light collection and focusing power - and the influence of the pump power on the laser rise time. Moreover, we mapped and identified the nature of the modes - FPM or WGM - involved in the lasing regime.
This work represent the first demonstration, of the potential of PINEM to study semiconductors population inversion dynamics, light absorption and field generated. In combination with the analytical capability of conventional TEM, PINEM enables to directly link the material’s atomic structure heterogeneity (e.g., defects, impurities, strain), the surrounding environment and the NWL geometry to the optical properties of the NWLs.
It must be stressed that the spatial, temporal, and energy resolutions of the UTEM used in this studied are respectively limited to 10 nm, 300 fs, and 1.0 eV, and we therefore expect the new generation of high-resolution instruments [schroder_laser-driven_2025, castioni_nanosecond_2025] to be able to dive even deeper in the physics of nanolasers or into more complex geometries, such as SPASERs [ellis_nanolasers_2024]. For example, monochromated electron microscopes [castioni_nanosecond_2025] may be able to resolve challenges related to mode hopping, cavity mode interference, and inter-mode dynamics.
4 Methods
TEM experiments requires an electron-transparent substrate. The NWLs were therefore deposited on a Holey carbon grid with an evaporated 100 nm Aluminium layer below the Carbon layer (the carbon layer playing the role of spacer between the Al layer and the NWL). This Al layer enables the determination of the pump-probe spatio-temporal overlap and seems to facilitate heat dissipation.
Electron energy detection scheme in the PINEM experiment : the electron energy spectrum were recorded using a Gatan PEELS 666 spectrometer with an ASI Cheetah 3 timepix detector (Fig. 1d).
Acknowledgements
The author thank Albert Polman and Nika Van Nielen for discussions and preliminary experiment. We also thank Mathieu Kociak for discussions and the all STEM group for their patience when the microscope was blocked at 200 keV. This work received support from the National Agency for Research under the program of future investment TEMPOS-CHROMATEM (ANR-10- EQPX-50). The authors acknowledge financial support from the CNRS-CEA “METSA” French network (FR CNRS 3507) on the platform LPS-STEM. The author acknowledge financial support of the French Agence Nationale de la Recherche (ANR), under the grant agreement ANR-23-CE09-0018 - LUTEM.
Author Contribution
C. Santini, with the assistance of S. Meuret, carried out the entire PINEM experiment and data analysis for the study, as well as building a portion of the PINEM setup required for the experiment. T.H. Ngo and B. Damilano created the samples with assistance from V. Brändl, S. Chenot, D. Lefebvre, and S. Vézian. L. Tizei carried out the EELS experiment and analyzed the data alongside S. Meuret. T. Fraysse and H. Lourenço-Martins assisted with the PINEM interpretation and data analysis. S. Weber and A. Teurtrie are in charge of the UTEM experiment and helped build a portion of the setup. S. Weber developped the acquisition software PyMoDAQ used in the experiment. A. Lloret oversees the optical setup for NWL characterization. C. Brimont and T. Guillet carried out the optical characterization and supervised the work with S. Meuret.